A beaker with 125 mL of an acetic acid buffer with a pH of 5.000 is sitting on a benchtop. The total molarity of acid and conjugate base in this buffer is 0.100 M. A student adds 6.80 mL of a 0.300 M HCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.740. Express your answer numerically to two decimal places. Use a minus ( ? ) sign if the pH has decreased.
Answer #1

Ans. Step 1: Using HH equation:   pH = pKa + log ([A-] / [AH])

            Where, AH = weak acid       ; A- = Conjugate base

Putting the values in above equation-

            5.000 = 4.740 + log ([A-] / [AH])

            Or, [A-] / [AH] = antilog (5.000 – 4.740) = 10-0.260

Or, [A-] = 0.5495 [AH]       - equation 1

Given, [AH] + [A-] = 0.100                        - equation 2

Comparing equation 1 and 2-

            [AH] + 0.5495 [AH] = 0.100

            Or, [AH] = 0.100 / 1.5495

            Hence, [AH] = 0.065

# Putting the value of [AH] in equation 2-

            [A-] = 0.100 – 0.065 = 0.035

# Step 2:

A = Weak acid A-= Conjugate base Ka of acid1.80E-05 Kb 10-14/ Kb 5.56E-10 pKa -log Ka 4.74 Original Buffer Strong Acid [AH] =

# Change in pH = final pH – Initial pH = 4.389 – 5.000 = -0.611